# Quantum Circuit Design using a Progressive Widening Enhanced Monte Carlo Tree Search

### v1

@@ Abstract
The performance of Variational Quantum Algorithms (VQAs) strongly depends on the choice of the parameterized quantum circuit to optimize. One of the biggest challenges in VQAs is designing quantum circuits tailored to the particular problem and to the quantum hardware. This article proposes a gradient-free Monte Carlo Tree Search (MCTS) technique to automate the process of quantum circuit design. It introduces a novel formulation of the action space based on a sampling scheme and a progressive widening technique to explore the space dynamically. When testing our MCTS approach on the domain of random quantum circuits, MCTS approximates unstructured circuits under different values of stabilizer Rényi entropy. It turns out that MCTS manages to approximate the benchmark quantum states independently from their degree of nonstabilizerness. Next, our technique exhibits robustness across various application domains, including quantum chemistry and systems of linear equations. Compared to previous MCTS research, our technique reduces the number of quantum circuit evaluations by a factor of 10 to 100 while achieving equal or better results. In addition, the resulting quantum circuits have up to three times fewer CNOT gates.
@@ 2 Related Work
Reinforcement learning is also a promising approach for QAS as shown by different frameworks developed for deep reinforcement learning agents [33, 34, 35, 36]. Those methods require a considerable amount of training data and computational resources. Moreover, deep reinforcement learning agents can easily get stuck in local minima as shown on the classical analogue NAS problems [37]. Recent advances in reinforcement learning also considered the effect of quantum noise in the QAS for the ground-state energy problem on three benchmark molecules [38].
@@
Inspired by [39], NAS problems have also been formulated in a tree representation to exploit Monte Carlo Tree Search (MCTS) techniques to design deep neural networks [40, 41]. Analogously, Meng et al. [17] proposed a tree representation of QAS and employed MCTS to design PQCs for quantum chemistry and condensed matter physics. Subsequently, Wang et. al. [18] have generalized the method, which achieves remarkable results on a wide range of applications, including the ground state energy problem for three different molecules, a system of linear equations, and more. The problem is formulated in a tree structure and a Nested MCTS (NMCTS) is employed for the search. Quantum circuits are sliced into a fixed number of empty layers and NMCTS fills them by choosing between a pre-defined pool of operations. Each layer of the quantum circuit corresponds to a tree level. The branching factor is given by the number of combinations that can be realized with the pool of operations. In this framework, an epoch is composed of the search technique acting on the topology of PQCs and a classical gradient-based technique acting on the angle parameters. This technique requires an expensive parameter optimization step at each epoch. Additionally, it requires fixing several hyperparameters that strongly affect the solution, such as the number of layers, which implicitly determines the number of gates in the PQC. However, these quantities are generally unknown and problem-dependent. Hence, an expensive hyperparameter tuning has to be performed on several quantities for any problem to find a suitable ansatz.
@@
This article addresses the key limitations encountered by MCTS approaches introducing a different tree-based action space formulation and some enhancements to the vanilla MCTS. We address the reader to Section 3.3 for the details. PWMCTS provides not only a higher level of automation but also significant improvements in the number of circuit evaluations required to design the PQC and its gates count (including CNOTs).
@@ 3 Progressive Widening enhanced Monte Carlo Tree Search
This section presents the Progressive Widening enhanced Monte Carlo Tree Search technique (PWMCTS) for the Quantum Architecture Search problem. In contrast to MCTS-based prior works [17, 18], we do not use gradient-based techniques during the search. It results in a gradient-free MCTS technique that designs both the topology and the parameters of PQCs. Note that, a gradient-based optimizer is employed for the standard fine-tuning of the parameters of the final circuit designed by MCTS.
The section is organized as follows. First, we formally define the QAS problem in Subsection 3.1. Then, we provide a general introduction to MCTS in Section 3.2. Finally, we present a detailed description of PWMCTS in Section 3.3.
@@ 3.1 Quantum Architecture Search
The $n$-qubit unitary operator $V$ consists of an ordered sequence of $L$ quantum gates $V_{i}$, picked from a certain universal gate set. In this notation, the initial state is |$\ket{0}=\ket{0}^{\otimes n}$ and $V_{i}$ is either a single-qubit or $2$-qubit operator tensored with the identity operator on the subspace of the unaffected qubit(s). If $V_{i}$ is a non-parameterized gate then $\theta_{i}$ is zero-dimensional. As a result, the vector of angle parameters is $\theta=(\theta_{0},\theta_{1},...,\theta_{l})\in[0,2\pi]^{l}$, with $l\leq L$. In the following, we will use the term variational quantum state |$\ket{\psi(\theta)}$ and quantum circuit interchangeably. Figure 1 represents the QAS problem on a generic PQC.
@@ 3.2 Monte Carlo Tree Search
The roll-out starts when the search reaches a leaf node. Hence, actions are selected in self-play until a terminal state is reached. This task might consist of playing plain random actions or – better – semi-random actions chosen according to a simulation strategy. In our agnostic approach, we work with a random roll-out. Simulation strategies based on heuristics have the potential to improve the performance of MCTS significantly [48].
@@ 3.3 Monte Carlo Tree Search for Quantum Architecture Search
We propose a Progressive Widening enhanced MCTS (PWMCTS) placed in a framework where each node corresponds to an $n$-qubit quantum circuit and each action corresponds to a particular circuit modification. The reward for PWMCTS is a real-valued function $R$ on the domain of $n$-qubit quantum circuits reaching its maximum value for the circuit’s optimal configuration. The objective function $C$ introduced in Section 3.1 is mapped onto the reward function through a problem-specific function $f$
@@
The initial circuit saved in the root node can be user-defined. By default, we start from the non-classical state generated by the quantum circuit composed by a Hadamard gate applied on each qubit, all initialized to |$\ket{0}$. PWMCTS explores the search space by sampling from four classes of allowed actions:
@@
    deleting from the circuit a gate at random position (D);
@@
Preliminary experiments on the value of the branching factor of the tree show that it highly impacts the performance of MCTS in our QAS formulation, see Figure 10. For this purpose, we equipped MCTS with a progressive widening technique [20]. It adaptively chooses the branching factor $k_{s}$ for each node, encoding the state $s$, in relation to the number of times the node gets visited $N_{s}$, according to the following formula
@@
Because we are interested in the whole sequence of gates an action-by-action search is employed to distribute the search time along all the levels of the tree [50, 51]. Once a level of the tree is sufficiently explored, PWMCTS commits to the action that has the best score so far. Specifically, the PWMCTS agent commits to a node once its number of visits reaches a certain percentage, $\rho$, of the total number of Monte Carlo iterations, $I$. Then, the search continues starting from the node it committed to. Note that $\rho$ is a hyperparameter that affects the depth of the tree and consequently the number of gates in the quantum circuit.
@@
Table 1 gives an overview of the hyperparameters of PWMCTS and the values used in the experiments. Note that, a further hyperparameter exploration may improve the results, especially on the progressive widening and action probability hyperparameters.
@@ 4.1 Variational Quantum Eigensolver
The Variational Quantum Eigensolver (VQE) is a NISQ approach to the ground state energy problem [5]. The objective is to minimize the cost function $C$ defined by the expectation value of the Hamiltonian $H$ of the system on the variational state |$\ket{\psi}=\ket{\psi(\theta)}$:
@@ 4.1.1 Experimental Results
The set of Hamiltonians used for all the experiments in this section have been generated from the quantum chemistry module in PennyLane [54] together with the electronic structure package PySCF. The Hamiltonian $H_{\mu}$ of a molecule $\mu$ is defined by fixing the atoms that compose it with their respective positions in the $3$-dimensional Cartesian coordinate system, the number of active electrons, and the number of active orbitals. All the numerical results are provided in atomic units (au) for the lengths and Hartree (Ha) for the energies. We execute the algorithm $10$ times for each experiment to collect statistics on independent events.
The results are compared with two classical benchmarks, the Self-Consistent Field (SCF) and the Full Configuration Interaction (FCI) methods. The SCF method is a mean-field approach that approximates the ground state using a single Slater determinant. It is a common computationally efficient benchmark. The FCI method provides an exact solution to the electronic problem within a given basis set. It captures all possible electronic correlations, serving as the standard for benchmarking quantum chemistry methods. However, it comes with a factorial-scaling computational cost. We refer the reader to Section 5.2 for a comparison with related quantum approaches.
@@
Figure 3 shows the numerical results of PWMCTS on the ground-state problem for H2. It converges to an approximate solution with only $1000$ iterations and less than $200$ iterations with the Adam Optimizer on the angle parameters. Increasing the value of $I$, PWMCTS designs PQCs that require fewer steps of the classical optimizer to fine-tune the parameters. Note that it does not obtain a result as good as the FCI method, but improves upon the SCF method. Compared to related quantum approaches it reaches the same local minima using considerably fewer quantum circuit evaluations, see Table 2.
@@
Figure 4 shows the numerical results of PWMCTS on the ground-state problem for H2O. Fixing $I\geq 5000$, PWMCTS achieves lower energies than the SCF method even without fine-tuning the parameters. After the fine-tuning, it converges consistently to the same value.
@@ 4.2.1 Experimental Results
Figure 8 shows the quantum ansatz provided by PWMCTS for the system of linear equation defined in Equations 16 and 17.
@@ 4.3 Quantum Oracle Approximation
In this application, we test PWMCTS in an unstructured problem setting. A dataset of quantum circuits randomly generated defines the target solutions. The motivation for this application is twofold. On one hand, we test our approach on a problem that is free from underlying symmetries of particular domains. On the other hand, we analyze PWMCTS as a technique to design quantum circuits that are difficult to simulate on a classical computer.
The classical simulability of quantum states is fundamental to assessing achievable quantum advantage. In this context, nonstabilizerness, commonly known as quantum magic, is a measure for quantum states that describes their difficulty to be characterized and manipulated on a classical computer [13, 57, 58, 59]. For this reason, the target random quantum circuits, are ranked according to their level of relevance based on their value of stabilizer Rényi entropy [22], chosen to measure the magic of quantum states because of its computational properties.
@@
In the following, we describe the dataset generation. All quantum circuits are initialized to |$\ket{0}^{\otimes n}$. Subsequently, we apply quantum gates sampled from the universal gate set composed of the Clifford generators (the Hadamard gate $H$, the phase gate $I$, the controlled-NOT) plus the non-Clifford $T$ gate. The problem instances in this domain are defined by target quantum circuits that differ in the number of qubits, $n$, number of gates, $g$, and magic. We generate $10$ quantum circuits for each combination ($n$, $g$) and select the two of them characterized by the lowest and highest values of stabilizer Rényi entropy $M_{2}$. The number of qubits varies in the set {4, 6, 8}, the number of gates in {5, 10, 15, 20, 30} and the magic {easy, hard}. As a result, we have $30$ different target circuits labeled by $(n,g,M_{2})$, each defining a different problem for PWMCTS. Note that, when we refer to the hardest (easiest) quantum circuit generated, it is relative to the quantum circuits in our dataset, hence, it does not correspond to the overall hardest (easiest) quantum circuit to simulate classically given the qubits and gates constraints. The number of $T$ gates involved in creating a quantum state is directly related to its stabilizer Rényi entropy [22]. However, high total quantum gates $g$ in the target quantum circuits do not guarantee high magic values. This is a consequence of the random process employed in the dataset generation. Note that, the universal gate set used for the data generation is different from that used by PWMCTS. This choice has been made to avoid the undesired structural connection between the problem and the model, which could affect the fairness of the experiments.
@@ 4.3.1 Experimental Results
In our experiments, we consider the algorithm successful when it manages to approximate with $\epsilon=0.05$ the $4$-qubits circuits, $\epsilon=0.1$ the $6$-qubits circuits, $\epsilon=0.2$ the $8$-qubits circuits. The difficulty of the problem increases with number of qubits. The choice of the values of $\epsilon$ and the size of quantum circuits allows to keep low the number of PWMCTS iterations $I$ and consequently the computation time.
The performance of our technique improves by increasing the number of PWMCTS iterations in all experiments $(n,g,M_{2})$, demonstrating the capabilities of our approach in this domain. However, there are some cases where the algorithm cannot find sufficiently good approximations. This is the case of the target $4$-qubits quantum circuits with $30$ gates.
On one hand, the problem’s difficulty is related to the number of qubits and the value of $\epsilon$. On the other hand, the number of gates used to generate the target circuit does not directly relate to the difficulty of the problem, which rather depends on the number of quantum gates needed to approximate the solution using the universal gate set $G$ used by PWMCTS, defined in Equation 8. Furthermore, we found that the nonstabilizerness of the target quantum states does not affect the performance of our technique. However, the experiments are based on a limited number of target quantum circuit samples.
Investigating the possibility of guiding PWMCTS towards solutions with low or high magic might be interesting in many domains. Depending on the goal, it can be beneficial to work with stabilizer states or in many others to build up quantum solutions far from the classical ones.

### v2

@@ Abstract
The performance of Variational Quantum Algorithms (VQAs) strongly depends on the choice of the parameterized quantum circuit to optimize. One of the biggest challenges in VQAs is designing quantum circuits tailored to the particular problem. This article proposes a gradient-free Monte Carlo Tree Search (MCTS) technique to automate the process of quantum circuit design. Our proposed technique introduces a novel formulation of the action space based on a sampling scheme and a progressive widening technique to explore the space dynamically. When testing our MCTS approach on the domain of random quantum circuits, MCTS approximates unstructured circuits under different values of stabilizer Rényi entropy. It turns out that MCTS manages to approximate the benchmark quantum states independently from their degree of nonstabilizerness. Next, our technique exhibits robustness across various application domains, including quantum chemistry and systems of linear equations. Compared to previous MCTS research, our technique reduces the number of quantum circuit evaluations by a factor of 10 up to 100 while achieving equal or better results. In addition, the resulting quantum circuits exhibit up to three times fewer CNOT gates, which is important for implementation on noisy quantum hardware.
@@ 2 Related Work
Reinforcement learning is also a promising approach for QAS, as shown by different frameworks developed for deep reinforcement learning agents [33, 34, 35, 36]. Those methods require a considerable amount of training data and computational resources. Moreover, deep reinforcement learning agents can easily get stuck in local minima as shown on the classical analogue NAS problems [37]. Recent advances in reinforcement learning also considered the effect of quantum noise in the QAS for the ground-state energy problem on three benchmark molecules [38].
Training-free approaches to QAS are also worth mentioning for the reduced computational cost required to evaluate quantum circuits. He et al. [39] proposed a method based on the directed acyclic graph representation of quantum circuits, where circuits are evaluated based on the number of paths in their graph representation and their expressibility. On one hand, this problem-agnostic method is computationally efficient. On the other hand, additional problem-dependent metrics are required for effective applications to VQAs. Expressibility of quantum circuits has also been recently used to pretrain a self-supervised learning model, which, when combined with Bayesian optimization, provides a robust solution for QAS [40]. Motivated by the limitations of the NISQ devices, Situ et al. [41] managed to develop a distributed approach to QAS, which also employs a training-free evaluation methodology.
@@
Inspired by [42], NAS problems have also been formulated in a tree representation to exploit Monte Carlo Tree Search (MCTS) techniques to design deep neural networks [43, 44]. Analogously, Meng et al. [17] proposed a tree representation of QAS and employed MCTS to design PQCs for quantum chemistry and condensed matter physics. Subsequently, Wang et. al. [18] have generalized the method, which achieves remarkable results on a wide range of applications, including the ground state energy problem for three different molecules, a system of linear equations, and more. The problem is formulated in a tree structure and a Nested MCTS (NMCTS) is employed for the search. Quantum circuits are sliced into a fixed number of empty layers and NMCTS fills them by choosing between a pre-defined pool of operations. Each layer of the quantum circuit corresponds to a tree level. The branching factor is given by the number of combinations that can be realized with the pool of operations. In this framework, an epoch is composed of the search technique acting on the topology of PQCs and a classical gradient-based technique acting on the angle parameters. This technique requires an expensive parameter optimization step at each epoch. Additionally, it requires fixing several hyperparameters that strongly affect the solution, such as the number of layers, which implicitly determines the number of gates in the PQC. However, these quantities are generally unknown and problem dependent. Hence, an expensive hyperparameter tuning has to be performed on several quantities for any problem to find a suitable ansatz.
@@
This article addresses the key limitations encountered by MCTS approaches introducing a different tree-based action space formulation and some enhancements to the vanilla MCTS. We refer the reader to Section 3.3 for the details. Our proposed PWMCTS approach provides not only a higher level of automation but also significant improvements in the number of circuit evaluations required to design the PQC and its gate count (including CNOTs).
@@ 3 Progressive Widening enhanced Monte Carlo Tree Search
This section presents the Progressive Widening enhanced Monte Carlo Tree Search technique (PWMCTS) for the Quantum Architecture Search problem. In contrast to MCTS-based prior research [17, 18], we do not use gradient-based techniques during the search. This results in a gradient-free MCTS technique that designs both the topology and the parameters of the PQCs. Note that a gradient-based optimizer is employed for standard fine-tuning of the parameters of the final circuit designed by MCTS.
The section is organized as follows. First, we formally define the QAS problem in Subsection 3.1. Then, we provide a general introduction to MCTS in Subsection 3.2. Finally, we present a detailed description of PWMCTS in Subsection 3.3.
@@ 3.1 Quantum Architecture Search
The $n$-qubit unitary operator $V$ consists of an ordered sequence of $L$ quantum gates $V_{i}$, picked from a certain universal gate set. In this notation, the initial state is |$\ket{0}=\ket{0}^{\otimes n}$ and $V_{i}$ is either a single-qubit or $2$-qubit operator tensored with the identity operator on the subspace of the unaffected qubit(s). If $V_{i}$ is a non-parameterized gate, then $\theta_{i}$ is zero-dimensional. As a result, the vector of angle parameters is $\theta=(\theta_{0},\theta_{1},...,\theta_{l})\in[0,2\pi]^{l}$, with $l\leq L$. In the following, we will use the term variational quantum state |$\ket{\psi(\theta)}$ and quantum circuit interchangeably. Figure 1 represents the QAS problem on a generic PQC.
@@ 3.2 Monte Carlo Tree Search
The roll-out starts when the search reaches a leaf node. Hence, actions are selected in self-play until a terminal state is reached. This task might consist of playing plain random actions or, better, semi-random actions chosen according to a simulation strategy. In our agnostic approach, we work with a random roll-out. Simulation strategies based on heuristics have the potential to improve the performance of MCTS significantly [51].
@@ 3.3 Monte Carlo Tree Search for Quantum Architecture Search
We propose a Progressive Widening enhanced MCTS (PWMCTS) placed in a framework where each node corresponds to an $n$-qubit quantum circuit and each action corresponds to a particular circuit modification. The reward for PWMCTS is a real-valued function $R$ in the domain of $n$-qubit quantum circuits reaching its maximum value for the circuit’s optimal configuration. The objective function $C$ introduced in Section 3.1 is mapped onto the reward function through a problem-specific function $f$
@@
The initial circuit saved in the root node can be user-defined. By default, we start from the non-classical state generated by the quantum circuit composed by a Hadamard gate applied to each qubit, all initialized to |$\ket{0}$. PWMCTS explores the search space by sampling from four classes of allowed actions:
@@
    deleting from the circuit a gate at a random position (D);
@@
Preliminary experiments on the value of the branching factor of the tree show that it highly impacts the performance of MCTS in our QAS formulation, see Figure 10. For this purpose, we equipped MCTS with a progressive widening technique [53, 20]. It adaptively chooses the branching factor $k_{s}$ for each node, encoding the state $s$, in relation to the number of times the node gets visited $N_{s}$, according to the following formula
@@
Because we are interested in the whole sequence of gates an action-by-action search is employed to distribute the search time along all levels of the tree [54, 55]. Once a level of the tree is sufficiently explored, PWMCTS commits to the action that has the best score so far. Specifically, the PWMCTS agent commits to a node once its number of visits reaches a certain percentage, $\rho$, of the total number of Monte Carlo iterations, $I$. Then, the search continues starting from the node it committed to. Note that $\rho$ is a hyperparameter that affects the depth of the tree and consequently the number of gates in the quantum circuit.
@@
Table 1 gives an overview of the hyperparameters of PWMCTS and the values used in the experiments. The values of the single configuration of hyperparameters used in in this article have been determined using a grid search around the reference values found in the literature [18, 19, 20, 59]. Note that, a further hyperparameter exploration may improve the results, especially on the progressive widening and action probability hyperparameters.
@@ 4.1 Variational Quantum Eigensolver
The Variational Quantum Eigensolver (VQE) is a NISQ approach to the ground-state energy problem [5]. The objective is to minimize the cost function $C$ defined by the expectation value of the Hamiltonian $H$ of the system on the variational state |$\ket{\psi}=\ket{\psi(\theta)}$:
@@ 4.1.1 Experimental Results
The set of Hamiltonians used for all the experiments in this section have been generated from the quantum chemistry module in PennyLane [58] together with the electronic structure package PySCF. The Hamiltonian $H_{\mu}$ of a molecule $\mu$ is defined by fixing the atoms that compose it with their respective positions in the $3$-dimensional Cartesian coordinate system, the number of active electrons, and the number of active orbitals. All the numerical results are provided in atomic units (au) for the lengths and Hartree (Ha) for the energies. We executed the algorithm $10$ times for each experiment to collect statistics on independent events.
The results are compared with two classical benchmarks, the Self-Consistent Field (SCF) and the Full Configuration Interaction (FCI) methods. The SCF method is a mean-field approach that approximates the ground state using a single Slater determinant. It is a common, computationally efficient benchmark. The FCI method provides an exact solution to the electronic problem within a given basis set. It captures all possible electronic correlations, serving as the standard for benchmarking quantum chemistry methods. However, it comes with a factorial-scaling computational cost. We refer the reader to Section 5.3 for a comparison with related quantum approaches.
@@
Figure 3 shows the numerical results of PWMCTS on the ground-state problem for H2. It converges to an approximate solution with only $1000$ iterations and less than $200$ iterations with the Adam Optimizer on the angle parameters. Increasing the value of $I$, PWMCTS designs PQCs that require fewer steps of the classical optimizer to fine-tune the parameters. Note that it does not obtain a result as good as the FCI method, but it improves upon the SCF method. Compared to related quantum approaches it reaches the same local minima using considerably fewer quantum circuit evaluations, see Table 2.
@@
Figure 4 shows the numerical results of PWMCTS on the ground-state problem for H2O. Fixing $I\geq 5000$, PWMCTS achieves lower energies than the SCF method even without fine-tuning the parameters. After fine-tuning, it converges consistently to the same value.
@@ 4.2.1 Experimental Results
Figure 8 shows the quantum ansatz provided by PWMCTS for the system of linear equations defined in Equations 16 and 17.
@@ 4.3 Quantum Oracle Approximation
In this application, we test PWMCTS in an unstructured problem setting. A dataset of quantum circuits randomly generated defines the target solutions. The motivation for this application is twofold. On the one hand, we test our approach on a problem that is free from the underlying symmetries of particular domains. On the other hand, we analyze PWMCTS as a technique to design quantum circuits that are difficult to simulate on a classical computer.
The classical simulability of quantum states is fundamental for assessing the achievable quantum advantage. In this context, nonstabilizerness, commonly known as quantum magic, is a measure for quantum states that describes their difficulty to be characterized and manipulated on a classical computer [13, 62, 63, 64]. For this reason, the target random quantum circuits are ranked according to their level of relevance based on their value of stabilizer Rényi entropy [22], chosen to measure the magic of quantum states due to its computational properties.
@@
In the following, we describe the dataset generation. All quantum circuits are initialized to |$\ket{0}^{\otimes n}$. Subsequently, we apply quantum gates sampled from the universal gate set composed of the Clifford generators (the Hadamard gate $H$, the phase gate $I$, the controlled-NOT) plus the non-Clifford $T$ gate. The problem instances in this domain are defined by target quantum circuits that differ in the number of qubits, $n$, number of gates, $g$, and magic. We generate $10$ quantum circuits for each combination ($n$, $g$) and select the two of them characterized by the lowest and highest values of stabilizer Rényi entropy $M_{2}$. The number of qubits varies in the set {4, 6, 8}, the number of gates in {5, 10, 15, 20, 30} and the magic {easy, hard}. As a result, we have $30$ different target circuits labeled by $(n,g,M_{2})$, each defining a different problem for PWMCTS. Note that, when we refer to the hardest (easiest) quantum circuit generated, it is relative to the quantum circuits in our dataset, hence, it does not correspond to the overall hardest (easiest) quantum circuit to simulate classically given the qubits and gates constraints. The number of $T$ gates involved in creating a quantum state is directly related to its stabilizer Rényi entropy [22]. However, high total quantum gates $g$ in the target quantum circuits do not guarantee high magic values. This is a consequence of the random process employed in the dataset generation. Note that, the universal gate set used for the data generation is different from that used by PWMCTS. This choice was made to avoid the undesired structural connection between the problem and the model, which could affect the fairness of the experiments.
@@ 4.3.1 Experimental Results
In our experiments, we consider the algorithm successful when it manages to approximate with $\epsilon=0.05$ the $4$-qubits circuits, $\epsilon=0.1$ the $6$-qubits circuits, $\epsilon=0.2$ the $8$-qubits circuits. The difficulty of the problem increases with the number of qubits. The choice of the values of $\epsilon$ and the size of quantum circuits allows to keep the number of PWMCTS iterations $I$ low and consequently the computation time.
The performance of our technique improves by increasing the number of PWMCTS iterations in all experiments $(n,g,M_{2})$, demonstrating the capabilities of our approach in this domain. However, there are some cases where the algorithm cannot find good approximations. This is the case of the target $4$-qubits quantum circuits with $30$ gates.
On the one hand, the problem’s difficulty is related to the number of qubits and the value of $\epsilon$. On the other hand, the number of gates used to generate the target circuit does not directly relate to the difficulty of the problem, which rather depends on the number of quantum gates needed to approximate the solution using the universal gate set $G$ used by PWMCTS, defined in Equation 8. Furthermore, we found that the nonstabilizerness of the target quantum states does not affect the performance of our technique. However, the experiments are based on a limited number of target quantum circuit samples.
Investigating the possibility of guiding PWMCTS towards solutions with low or high magic might be interesting in many domains. Depending on the goal, it can be beneficial to work with stabilizer states or to build up quantum solutions far from the classical ones.